OpenStudy (anonymous):

if 49=(7y)2 then y could equal to a.-1 b.1 c.7. a,b,c??

OpenStudy (john_es):

Answer (a) and (b) are correct.

OpenStudy (anonymous):

ecxplain ..please

OpenStudy (john_es):

\[49=(7y)^2\Rightarrow {49}=49y^2\Rightarrow \frac{49}{49}=y^2\Rightarrow 1=y^2\Rightarrow \pm\sqrt{1}=\sqrt{y^2}\Rightarrow y=\pm1\]